OpenStudy (anonymous):

How do you do this? in point slope form? (-8,3) m + - 3/5?

OpenStudy (icalibear):

point slope form = y = mx + b

OpenStudy (anonymous):

I need the answer i was absent and it says write in point slope form of n equation for each line passing through the given point and having the given slope.

OpenStudy (anonymous):

the format is y-y1=m(X-X1)

OpenStudy (icalibear):

y - 3 = m(x + 8)